In excersise 33-41,Use the given information to find the given amount.
40. Food bill: $20
Sales tax: 5%
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

20+.05*20=20+1=$21 THE FINAL PRICE WITH TAX.